# Precovery Observations of 3I/ATLAS from TESS Suggests Possible Distant Activity

### v1

@@ 5 Data and Software Availability
The raw TESS FFIs are available for bulk download on the Mikulski Archive for Space Telescopes.444https://archive.stsci.edu/tess/bulk_downloads/bulk_downloads_ffi-tp-lc-dv.html All of the analysis Python scripts, data behind the figures, and Python scripts for reproducing the figures presented here are available on GitHub.555https://github.com/afeinstein20/atlas-tess Additional, larger, data products will be made available on Zenodo upon publication.
@@ 6 Acknowledgments
We thank Henry Hsieh, Davide Farnocchia, and Marco Micheli for thoughtful insights. A.D.F. acknowledges funding from NASA through the NASA Hubble Fellowship grant HST-HF2-51530.001-A awarded by STScI. D.Z.S. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303553. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time.

### v2

@@ 5 Data and Software Availability
The calibrated TESS FFIs are available for bulk download on the Mikulski Archive for Space Telescopes.444https://archive.stsci.edu/tess/bulk_downloads/bulk_downloads_ffi-tp-lc-dv.html All of the analysis Python scripts, data behind the figures, and Python scripts for reproducing the figures presented here are available on GitHub.555https://github.com/afeinstein20/atlas-tess Additional, larger, data products will be made available on Zenodo upon publication.
@@ 6 Acknowledgments
We thank Henry Hsieh, Davide Farnocchia, and Marco Micheli for thoughtful insights. We thank the anonymous referee for their comments which have improved the quality and clarity of this manuscript. A.D.F. acknowledges funding from NASA through the NASA Hubble Fellowship grant HST-HF2-51530.001-A awarded by STScI. D.Z.S. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303553. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time.
@@ 7 Author Contributions
Appendix A Background Removal
In Figure 10 we present the models and residuals tested to remove the background properties in the TESS FFIs. These experiments demonstrated that the Savitsky-Golay filter (with a window length of 307) presents the optimal fit to the observations without overfitting the faint signal of 3I/ATLAS for Camera 2 CCD 3. The larger window length is also able to recover the target but improperly fits the sharp decline of the orbital ramp. The 2nd-order polynomial does not fit the Camera 1 CCD 2 data well due to the strong systematics present in the observations. Similarly to Camera 2 CCD 3, the second image is best fit by the Savitsky-Golay filter with a window length of 307. Therefore, we select this background to subtract from the calibrated FFIs in the entirety of the analysis presented in this work.
Appendix B Aperture Selection
The TESS pixel response function (PRF) can change substantially over a given camera’s FOV. Additionally, the PRF is slightly chromatic and can vary with temperature. The TESS PRF model assumes that 24% of the total source flux will be contained in the central pixel if the target is perfectly centered. If the target is offset, then the total amount of flux decreases. It is particularly important to take this into account when trying to extract the flux from a moving target. To combat these effects, larger apertures are more robust against these flux losses. However, implementing such large apertures is also particularly challenging for crowded fields because they may include flux contributions from nearby sources.
To investigate the efficacy of both competing effects, we perform a set of experiments on a variety of apertures to extract the light curve of 3I/ATLAS. The subset of apertures investigated here are provided in A. D. Feinstein et al. (2019), ranging from a single pixel to a area. The results of these tests are presented in Figure 11 and Figure 12. These figures show the full light curve as well as the light curve binned to 36 minutes. We also run the binned light curve through a Lomb-Scargle periodogram to search for potential evidence of a rotation period.
Figure 11 demonstrates that the light curve is dominated by noise at hours. Additionally, there is no strong evidence of periodicity for hours. While there are stronger peaks at hours, the majority of these periodograms have peaks at identical times within this time range. We note that the periodograms shown in rows 5-8 of Figure 11 contain similar peaks at hours, which could be consistent with the reported hours (R. de la Fuente Marcos et al., 2025). However, because the same signal is not seen across all extracted light curves, we attribute it to instrumental systematics.
We repeat this same test with the data extracted from Camera 1 CCD 2. These results are presented in Figure 12. Similarly to the previous analysis of the other TESS image, the hours is dominated by noise. Contrary to the previous analysis however, there is a strong peak between hours. Additionally, we find a consistent strong double-peaked feature at hours. The 16 hour peak is consistent with recovered by (R. de la Fuente Marcos et al., 2025; T. Santana-Ros et al., 2025). These signals are strong in every extracted light curve. None of the stronger peaks seen in the light curves extracted from Camera 1 CCD 2 are consistent with those from Camera 2 CCD 3. Additionally, because there are multiple strong peaks and no observable photometric variability in all data, it is challenging to determine whether the signals are instrumental or astrophysical in origin. Nevertheless, because the periodograms between TESS data sets are inconsistent, we conclude that we are unable to measure a rotation period from these observations.